OpenStudy (anonymous):

(square root 2)tan x cos x - tan x=0 solve for x I get that I will need to square the whole problem but I'm not sure what it will look like after squaring

OpenStudy (phi):

can you re-write the question? It is not clear what is under the square root.

OpenStudy (anonymous):

yes \[\sqrt{2} \tan x \cos x-\tan x=0\]

OpenStudy (anonymous):

so would it be \[2 \tan ^{2} x \cos^{2} x-\tan^{2} x=0\] or is that wrong?

OpenStudy (phi):

Is the original this \[ \sqrt{2 \tan x \cos x} - \tan x = 0 \]?

OpenStudy (anonymous):

no the square root is only over the 2

OpenStudy (phi):

ok, then you do not square anything. factor out tan x what do you get ?

OpenStudy (anonymous):

\[\tan x (\sqrt{2} \cos x-1)=0\] then do I set both equal to zero and solve or not yet?

OpenStudy (phi):

then do I set both equal to zero. yes. this is the same old idea: if you multiply two terms like A * B =0 to get 0, then the obvious answer is either A is 0 or B is 0 (or both at the same time) in this case you have tan x =0 or (sqrt(2) cos(x) -1) =0

OpenStudy (anonymous):

thank you so much for your patience!